OpenStudy (smart8thgrader):

Which term could you combine with x²? Question 3 options: 2x x/2 3x² 2/x

OpenStudy (smart8thgrader):

please help me I dont understand

OpenStudy (slickster256):

I believe its c because they have a common variable

Nnesha (nnesha):

if the variable is same with same exponent then you can combine their coefficients here is an example 2x+3x = (3+2)x =5x 2 and 3 both have the same variable same exponent so i can combine them

Nnesha (nnesha):

2x +3x^2 can you combine them ?

OpenStudy (smart8thgrader):

I think

OpenStudy (smart8thgrader):

Is it c?

Nnesha (nnesha):

no you `can't` bec exponents aren't the same \[\huge\rm 2x^1 +3x^2\]

Nnesha (nnesha):

what about this one \[\huge\rm 2x^3 +3y^3\] can you combine 2 and 3 ?

OpenStudy (smart8thgrader):

IDK

Nnesha (nnesha):

well if the `variables` and the `exponents` are the same then you can combine their coefficients :=)

Nnesha (nnesha):

\[\huge\rm 2x^3 +3y^3\]now tell me can you combine them ?

OpenStudy (smart8thgrader):

Which is equivalent to 3x + 4y + 3x + 4y? Question 4 options: 6x + 8y 2(3x + 4y) Answers I & II None of the above

Vocaloid (vocaloid):

3x + 3x = ?

Nnesha (nnesha):

same question `combine like terms`

Nnesha (nnesha):

if you understand the definition you will be able t solve rest of the questions